OpenStudy (anonymous):

find the are of a surface generated when the curve y=1/2x^3 + 1/2x^-1 from x=1 to x=3 is rotated about the x-axis.

OpenStudy (anonymous):

Solid of revol problem. Washer or shell?

OpenStudy (anonymous):

is with surface formula s=2pi int f(x) sqroot 91+(f(x))^2) dx

OpenStudy (anonymous):

sorry no 9.. (1+(f(x))^2)

OpenStudy (anonymous):

ohh thats shell ok cool.. i started but i am stuck...

OpenStudy (anonymous):

This should be easy because we have no intersections to worry about \[2\pi \int _a^bf(x) \sqrt{ f'(x)^2+1}dx\]

OpenStudy (anonymous):

Find F'(x)

OpenStudy (anonymous):

yea i have f'(x)=1/2x^2 - 1/2x^-2

OpenStudy (anonymous):

\[F(x)=\frac{1}{2}x^2+\frac{1}{2}x^{-1}\] wait is it the function?

OpenStudy (anonymous):

the ftc is 1/6x^3 + 1/2x^-1

OpenStudy (anonymous):

so f'(x)=1/2x^2 - 1/2x^-2

OpenStudy (anonymous):

square that

OpenStudy (anonymous):

am i rigth?

OpenStudy (anonymous):

ok i got 1/4x^4 - 1/2 + 1/4x^-4

OpenStudy (anonymous):

then i add 1

OpenStudy (anonymous):

yes what do you have after adding one?

OpenStudy (anonymous):

same different sign

OpenStudy (anonymous):

(1/2x^2 + 1/2x^-2)^2

OpenStudy (anonymous):

** different at the negative sign**

OpenStudy (anonymous):

F'(x)^2=\[\left(-\frac{1}{2 x^2}+\frac{x^2}{2}\right)^2\] 1+f'(x)^2 \[1+\left(-\frac{1}{2 x^2}+\frac{x^2}{2}\right)^2\] some algebra later \[\frac{1+2 x^4+x^8}{4 x^4}\]

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

1/2x^2 + 1/2x ^-2

OpenStudy (anonymous):

second term power negative

OpenStudy (anonymous):

so is in the denominator

OpenStudy (anonymous):

yes it is in denominator

OpenStudy (anonymous):

u didnt do it like that y put it in the nominator'

OpenStudy (anonymous):

\[F=\frac{1}{6}x^3+\frac{1}{2}x^{-1}\] \[F'=\frac{1}{2}x^2-\frac{1}{2}x^{-2}\] Rewritten as \[F'=\frac{1}{2}x^2-\frac{1}{2x^2}\]

OpenStudy (anonymous):

\[F'^2=\left(-\frac{1}{2 x^2}+\frac{x^2}{2}\right)^2\] \[F'^2+1=1+\left(-\frac{1}{2 x^2}+\frac{x^2}{2}\right)^2\] Rewritten as \[\frac{1+2 x^4+x^8}{4 x^4}\] which can be written as perfect square \[\frac{\left(x^4+1\right)^2}{\left(2x^2\right)^2}\] Since we are taking square root, we lose square \[\frac{\left(x^4+1\right)}{\left(2x^2\right)}\]

OpenStudy (anonymous):

\[2\pi \int _a^b\left(\frac{1}{6}x^3+\frac{1}{2}x^{-1}\right)\frac{\left(x^4+1\right)}{\left(2x^2\right)}dx\] \[2\pi \int _a^b\frac{1}{4 x^3}+\frac{x}{3}+\frac{x^5}{12}dx\] Simple integral from here

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

i just finish the problem.. i had 209/9 pi sq units...lol